OpenStudy (anonymous):

A student raises her grade average from a 75 to 90 what was the percent of increase in the students grade round your answers to the nearest tenth of a percent if NECESSARY A. 20% B. 16.7% C. 8.3% D. 15%

OpenStudy (anonymous):

@Brooklyn_Boiii @Hero @kxttxn @Missiey @ParthKohli @Pawanyadav

OpenStudy (anonymous):

look im bad a percentages but i think it's A

OpenStudy (anonymous):

@phi

OpenStudy (phi):

first, write the fraction: amount of change divided by original amount can you do that ?

OpenStudy (anonymous):

um not really

OpenStudy (phi):

what is the change (difference) between 75 and 90 ?

OpenStudy (anonymous):

75 is a smaller amount then 90

OpenStudy (phi):

yes, but by how much ?

OpenStudy (anonymous):

15

OpenStudy (phi):

yes, the change from 75 to 90 is 15 15 goes in the top of the fraction. The bottom of the fraction is the original amount. What is the original number?

OpenStudy (phi):

the original number is 75 (and then it changes to 90) to find the percent change we write the fraction \[ \frac{15}{75}\]

OpenStudy (phi):

Can you simplify that fraction and change it into a decimal ?

OpenStudy (anonymous):

15/75 reduces to 3/25 which is 12% then as a decimal = 0.12 @phi

OpenStudy (phi):

that is the right idea, but your arithmetic is wrong you can divide top and bottom by 3 to get 5/25 then divide top and bottom by 5 to get 1/5 that is 0.2 or 20%

OpenStudy (anonymous):

so i was right it's 20%? @phi

OpenStudy (anonymous):

or 15%

OpenStudy (phi):

Being "right" means being able to do the problem (not just the final answer) If you are running a race and get to the finish line, it does not count if you used a car. In other words, you really want to know *how to get the 20%* So try hard to learn how to do the problem.

OpenStudy (phi):

simplify 15/75 (you do not get 3/25)

OpenStudy (anonymous):

wait ur confusing me what is it 20 or 15 @phi

OpenStudy (phi):

can you simplify 15/75 ?

OpenStudy (anonymous):

=0.2 @phi

OpenStudy (anonymous):

@tom982 may u tell me

OpenStudy (anonymous):

or u @maddie82502 idk if it's 15 or 20

OpenStudy (phi):

yes 0.2 and that is 20%

OpenStudy (anonymous):

Yes \(\frac{15}{75}=\frac{1}{5}=0.2\) which, as a percentage, is \(20\)%.

OpenStudy (maddie82502):

I think its 20%

OpenStudy (anonymous):

kk thanks guy @maddie i want to fan everybody so can u fan @phi and i fan u and @phi fans @tom982

OpenStudy (anonymous):

Oh don't worry about that, I'm not fussed about medals/fans.

OpenStudy (anonymous):

oh ok well can u fan me then @maddie82502 idk

OpenStudy (maddie82502):

I fanned everyone

OpenStudy (anonymous):

oh ok thanks didnt know u can fan everyone

OpenStudy (anonymous):

oh wait im so sorry i ment medals

OpenStudy (anonymous):

i medaled u @maddie82502 u should of did @phi and i wont be medaled

OpenStudy (maddie82502):

im srry

OpenStudy (anonymous):

it's ok